# Bridging Research Gaps Between Academic Research and Legal Investigations of Algorithmic Discrimination

### v1

@@ Introduction
The legal framework for addressing algorithmic discrimination in the US is built upon the 5th and 14th Amendments to the U.S. Constitution, which guarantee fundamental rights related to due process and equal protection under the law to all persons, and the numerous civil rights laws that reinforce and extend these protections, including the Civil Rights Act of 1964 and Fair Housing Act of 1968, which create avenues for state and private action. Although the use of artificial intelligence (AI) technology is relatively new, courts and regulators have applied existing authorities to emerging algorithmic systems and resulting court decisions (such as (Mobley 2024)). States and localities have also started to enact new laws to regulate AI, such as the Colorado Artificial Intelligence Act of 2024, which creates a duty of reasonable care to guard against known or reasonably foreseeable risks of algorithmic discrimination  (Co. General Assembly 2024). Illinois’ HB3773 protects employees from AI-related discrimination in employment contexts (Ill. General Assembly 2023), following on city-level efforts such as New York City’s Local Law 144 (Wright et al. 2024). Such laws have been motivated in part by reports across multiple sectors demonstrating how automated decision-making can perpetuate discrimination through, e.g., hiring algorithms that systematically downgrade qualified female candidates (Dastin 2018), tenant screening systems that disproportionately reject minority applicants (Roth 2024), and healthcare allocation algorithms that provide lower levels of rehabilitation care to elderly patients (Ross and Herman 2023). Automated decision-making can perpetuate or amplify historical patterns of discrimination, even without explicit intent.
@@
Although disparate treatment and disparate impact provide the main legal frameworks for analyzing algorithmic discrimination, other relevant legal concepts are worth noting. One example is the four-fifths rule (also known as the 4/5 rule or 80% rule). This general rule of thumb, as promulgated by the Guidelines of the US Equal Employment Opportunity Commission (EEOC), is a method for determining whether the selection rate for one group is “substantially” different than the selection rate of another group, potentially evincing disparate impact-based discrimination (EEOC 2023). The rule states that one rate is substantially different than another if their ratio is less than four-fifths (or 80%) and has been endorsed by the Supreme Court in Duke v. Griggs Power Co. (Griggs 1971). Despite being one component of a broader legal framework, and despite the EEOC itself indicating that the four-fifths rule is “merely a rule of thumb” that may be “inappropriate under certain circumstances,” it has been disproportionately emphasized in ML literature, often misrepresenting actual disparate impact doctrine (Watkins and Chen 2024).
@@
Algorithmic discrimination oversight in the US operates through multiple legal channels as shown in Figure 1. While shown separately, in practice, these pathways may intersect, for example, with the evaluation of a private right of action or class action becoming the basis for an agency enforcement action. Federal agencies like the Federal Trade Commission (FTC), EEOC, and Consumer Financial Protection Bureau (CFPB) have historically taken a leading role in enforcing anti-discrimination protections. In recent years, the FTC has taken action against Rite Aid’s facial recognition safeguards (Federal Trade Commission 2023), and been urged to initiate action against HireVue’s AI assessment tools (Electronic Privacy Information Center 2019), and Aon’s employment screening technology (American Civil Liberties Union Foundation 2024) while the EEOC has provided guidance on AI hiring tools and investigated automated discrimination (EEOC 2023). Beyond agency enforcement, private litigation through class action suits and individual claims (Figure 1) establish precedents for applying discrimination frameworks to algorithms, exemplified by Mobley v. Facebook and Mobley v. Workday (Mobley 2016, 2024). Civil society organizations drive change through strategic litigation, as seen in the settlement between National Fair Housing Alliance (NFHA) and Facebook (National Fair Housing Alliance 2018) and the American Civil Liberties Union’s credit scoring algorithm challenges (Sandvig 2016). Several of these actions are described in Table 1 and were analyzed as part of this project.
While federal and private actors shape enforcement, state and local governments have enacted their own algorithmic accountability legislation. New York City’s Local Law 144 requires employment decision tool audits (Groves et al. 2024), Illinois mandates AI interview disclosures (Ill. General Assembly 2020), and, starting in 2026, Colorado will require developers and deployers of certain AI systems to use ”reasonable care” to protect consumers from the risks of algorithmic discrimination  (Co. General Assembly 2024). However, similar efforts have faced resistance in other states (Bedayn 2024), and by the Trump White House  (White House 2025). Regulatory approaches, including agency rulemaking like the CFPB’s automated credit guidance (CFPB 2023), industry self-regulation, academic partnerships, and international laws such as the AI Act of the European Union (EU 2024) complement domestic law making and enforcement activities. Comprehensive surveys of regulatory measures emphasize the evolving nature of U.S. legal practices in addressing different types of algorithmic discrimination (Wang, Chen, and Liu 2024).
@@ Methods
Each enforcement action centers on allegations of algorithmic discrimination. These cases broadly fall into two categories: (1) actions involving algorithms alleged to explicitly use protected characteristics, such as Apple Card’s screening algorithm’s alleged gender discrimination (app 2021a) and Allstate’s credit underwriting algorithm’s alleged racial discrimination (DeHoyos 2007), or (2) cases in which algorithms demonstrated performance issues or inadequate safeguards, as exemplified by Rite Aid’s facial recognition system (FTC 2024).
@@
The cases have varied outcomes: several have reached settlements (Roth 2024; Milstein 2024; FTC 2024; DOJ 2024; DeHoyos 2007), some found no wrongdoing (app 2021b), and ongoing cases continue to shape evolving standards for algorithmic auditing and compliance (Huskey 2022; Estate of Gene B. Lokken et al. ; Mobley 2024).
The legal justifications and remediation approaches in these cases highlight critical areas where ML research can better support enforcement. Cases frequently rely on anti-discrimination frameworks like the Fair Housing Act and Equal Credit Opportunity Act (ECOA) to establish protected characteristics and determine permissible algorithmic inputs  (DOJ 2024; DeHoyos et al. 2007; Milstein 2024). However, enforcers face several challenges when attempting to quantify discrimination: while some cases successfully demonstrated disparate impact through statistical evidence (DeHoyos 2007), others struggled with protected group imputation and data access barriers (Ross and Herman 2023; DOJ 2024). The remediation approaches ordered through settlements reveal the practical need for research on algorithm modification techniques that maintain accuracy while reducing bias. This is particularly evident in cases where companies were required to redesign their algorithms (Milstein 2024), highlighting the gap between theoretical fairness improvements and deployable solutions. Financial penalties and technology bans, while effective enforcement tools, underscore the necessity of scrupulous algorithmic investigations (DOJ 2024; Milstein 2024; FTC 2024). These enforcement patterns inform our identified research opportunities, particularly around developing standardized disparate impact measurements and techniques for finding equally accurate, less discriminatory alternatives.
@@
Under US civil rights law, defendants in discrimination cases can justify practices with disparate impact if they demonstrate “business necessity.” However, this defense fails if plaintiffs can show the existence of less discriminatory alternatives that achieve similar business objectives. This concept has direct relevance to algorithmic discrimination cases. In 2022, a class action lawsuit of homeowners alleged that State Farm’s homeowner insurance claims processing algorithms subjected Black policyholders to greater scrutiny. The complaint specifically questioned, “Whether substantially equally or more valid alternative means of claim processing are available that would eliminate or reduce the discriminatory impact [of State Farm’s claims processing policy]” (Huskey 2022). As of writing, the investigation remains ongoing as the plaintiffs have demonstrated a probable likelihood that their allegations have merit (United States District Court 2023). In 2007, a similar case emerged when plaintiffs sued Allstate for allegedly charging racially discriminatory insurance premiums. While denying these claims, Allstate agreed to implement a new pricing algorithm as part of the settlement (DeHoyos et al. 2007). Though the settlement documents do not describe the development of this alternative algorithm, it seems reasonable to assume that it was designed to balance non-discrimination with business performance requirements. Legal scholars have established that defendants have an affirmative duty to search for less discriminatory alternatives in many contexts (Black et al. 2024), and recent work has begun operationalizing this search in fair lending contexts (Gillis, Meursault, and Ustun 2024).
@@
In 2021, the New York State Department of Financial Services (NYDFS) investigated Apple based on consumer complaints about the Apple Card and the potential for algorithmic discrimination. After a lengthy investigation, the report found no evidence of unlawful discrimination against applicants under fair lending law. Part of the investigation focused on individuals who sent discrimination complaints to the NYDFS: “In each instance, the Bank was able to identify the factors that led to the credit decisions, such as credit score, indebtedness, income, credit utilization, missed payments, and other credit history elements” (app 2021b). Using regression analysis, investigators were able to determine that the Apple Card lending policy—and the underlying statistical model—would not produce disparate impacts (app 2021b).
@@
Our analysis suggests several critical areas for future research in quantifying algorithmic disparate impact. As algorithmic systems become more complex, they require sophisticated measurement approaches that can: (1) isolate interaction effects between protected attributes and other variables, (2) track temporal drift in model behavior and fairness metrics, and (3) model complex downstream economic consequences. Machine learning concepts such as distribution shift, causal inference, and feature attribution are crucial for these research questions. However, more work is needed to develop practical solutions. For example, in lending discrimination, impact quantification must account for both immediate effects of loan denials and long-term consequences for wealth accumulation and inter-generational mobility. Methods from econometrics and causal inference may help address this change. While prior works advocate for causal frameworks in measuring disparate impact (Xiang and Raji 2019; Xiang 2020), the field lacks validated methodologies and case studies demonstrating how to compute these measures in ways that both withstand legal scrutiny and support specific damage calculations. The few examples that do exist rely on quasi-experimental methods that are often infeasible to implement with the available data (Arnold, Dobbie, and Hull 2021). Recent work has highlighted that regulatory approaches must consider threshold variations across different populations (Meursault, Gillis, and Ustun 2024) and has begun developing frameworks for explainable fairness in regulatory auditing contexts (O’Neil, Lehr, and Vaithianathan 2024). This methodological gap significantly hampers enforcement actions seeking concrete remedies for affected populations.
@@
Without accurate demographic information, investigators may struggle to establish statistical evidence of discriminatory patterns and face heightened burdens of proof in legal proceedings. The lack of reliable race and ethnicity data particularly hampers class certification efforts, as courts require clear methods for identifying affected individuals. As an extreme case, certain sectors may forbid the consideration—and therefore collection—of these groups, such as the Equal Credit Reporting Act, which prohibits creditors from considering information about protected attributes in any aspect of a credit transaction (CFPB 2011; reg 2021). This lack of recorded information affected the Allstate 2007 settlement regarding allegedly racially discriminatory insurance premiums because “Allstate does not maintain information on the race of its policyholders”, and therefore notice could not be given to prospective class members (DeHoyos et al. 2007).

### v2

@@ Introduction
The legal framework for addressing algorithmic discrimination in the US is built upon the 5th and 14th Amendments to the Constitution, which provide due process and equal protection safeguards against state action, and a body of federal civil rights laws extending these protections to private action. These laws, including the Civil Rights Act of 1964 and the Fair Housing Act of 1968, create mechanisms to challenge discrimination in areas like employment, housing, and public accommodations. Courts and regulators have applied existing authorities to emerging algorithmic systems, resulting in enforcement actions and court decisions (e.g. such as (Mobley 2024)). States and localities have also started to enact new laws to regulate AI, such as the Colorado Artificial Intelligence Act of 2024, which creates a duty of reasonable care to guard against known or reasonably foreseeable risks of algorithmic discrimination  (Co. General Assembly 2024). Illinois’ HB3773 protects employees from AI-related discrimination in employment contexts (Ill. General Assembly 2023), following on city-level efforts such as New York City’s Local Law 144 (Wright et al. 2024). Such laws have been motivated in part by reports across multiple sectors demonstrating how automated decision-making can perpetuate discrimination through, e.g., hiring algorithms that systematically downgrade qualified female candidates (Dastin 2018), tenant screening systems that disproportionately reject minority applicants (Roth 2024), and healthcare allocation algorithms that provide lower levels of rehabilitation care to elderly patients (Ross and Herman 2023). Automated decision-making can perpetuate or amplify historical patterns of discrimination, even without explicit intent.
@@
Although disparate treatment and disparate impact provide the main legal frameworks for analyzing algorithmic discrimination, other relevant legal concepts are worth noting. One example is the four-fifths rule (also known as the 4/5 rule or 80% rule). This general rule of thumb, as promulgated by the Guidelines of the US Equal Employment Opportunity Commission (EEOC), is a method for determining whether the selection rate for one group is “substantially” different than the selection rate of another group, potentially evincing disparate impact-based discrimination (EEOC 2023). The rule states that one rate is substantially different than another if their ratio is less than four-fifths (or 80%) and has been endorsed by the Supreme Court in Griggs v. Duke Power Co. (Griggs 1971). Despite being one component of a broader legal framework, and despite the EEOC itself indicating that the four-fifths rule is “merely a rule of thumb” that may be “inappropriate under certain circumstances,” it has been disproportionately emphasized in ML literature, often misrepresenting actual disparate impact doctrine (Watkins and Chen 2024).
@@
Algorithmic discrimination oversight in the US operates through multiple legal channels as shown in Figure 1. While shown separately, in practice, these pathways may intersect, for example, with the evaluation of a private right of action or class action becoming the basis for an agency enforcement action. Federal agencies like the Federal Trade Commission (FTC), EEOC, and Consumer Financial Protection Bureau (CFPB) have historically taken a leading role in enforcing anti-discrimination protections. In recent years, the FTC has taken action against Rite Aid’s facial recognition safeguards (FTC 2023), and been urged to initiate action against HireVue’s AI assessment tools (Electronic Privacy Information Center 2019), and Aon’s employment screening technology (American Civil Liberties Union Foundation 2024) while the EEOC has provided guidance on AI hiring tools and investigated automated discrimination (EEOC 2023). Beyond agency enforcement, private litigation through class action suits and individual claims (Figure 1) establish precedents for applying discrimination frameworks to algorithms, exemplified by Mobley v. Facebook and Mobley v. Workday (Mobley 2016, 2024). Civil society organizations drive change through strategic litigation, as seen in the settlement between National Fair Housing Alliance (NFHA) and Facebook (National Fair Housing Alliance 2018) and the American Civil Liberties Union’s credit scoring algorithm challenges (Sandvig 2016). Several of these actions are described in Table 1 and were analyzed as part of this project.
While federal and private actors shape enforcement, state and local governments have enacted their own algorithmic accountability legislation. New York City’s Local Law 144 requires employment decision tool audits (Groves et al. 2024), Illinois mandates AI interview disclosures (Ill. General Assembly 2020), and, starting in 2026, Colorado will require developers and deployers of certain AI systems to use ”reasonable care” to protect consumers from the risks of algorithmic discrimination  (Co. General Assembly 2024). However, similar efforts have faced resistance in other states (Bedayn 2024), and by the Trump White House  (White House 2025). Regulatory approaches, including agency rulemaking like the CFPB’s automated credit guidance (CFPB 2023), industry self-regulation, academic partnerships, and international laws such as the AI Act of the European Union (EU 2024) complement domestic law making and enforcement activities. Comprehensive surveys of regulatory measures emphasize the evolving nature of U.S. legal practices in addressing different types of algorithmic discrimination (Wang et al. 2024).
@@ Methods
Each enforcement action centers on allegations of algorithmic discrimination. These cases broadly fall into two categories: (1) actions involving algorithms alleged to explicitly use protected characteristics, such as Apple Card’s screening algorithm’s alleged gender discrimination (New York State Department of Financial Services 2021a) and Allstate’s credit underwriting algorithm’s alleged racial discrimination (DeHoyos 2007), or (2) cases in which algorithms demonstrated performance issues or inadequate safeguards, as exemplified by Rite Aid’s facial recognition system (FTC 2024).
@@
The cases have varied outcomes: several have reached settlements (Roth 2024; Milstein 2024; FTC 2024; DOJ 2024; DeHoyos 2007), some found no wrongdoing (New York State Department of Financial Services 2021b), and ongoing cases continue to shape evolving standards for algorithmic auditing and compliance (Huskey 2022; Estate of Gene B. Lokken et al. 2024; Mobley 2024).
The legal justifications and remediation approaches in these cases highlight critical areas where ML research can better support enforcement. Cases frequently rely on anti-discrimination frameworks like the Fair Housing Act and Equal Credit Opportunity Act (ECOA) to establish protected characteristics and determine permissible algorithmic inputs  (DOJ 2024; DeHoyos et al. 2001; Milstein 2024). However, enforcers face several challenges when attempting to quantify discrimination: while some cases successfully demonstrated disparate impact through statistical evidence (DeHoyos 2007), others struggled with protected group imputation and data access barriers (Ross and Herman 2023; DOJ 2024). The remediation approaches ordered through settlements reveal the practical need for research on algorithm modification techniques that maintain accuracy while reducing bias. This is particularly evident in cases where companies were required to redesign their algorithms (Milstein 2024), highlighting the gap between theoretical fairness improvements and deployable solutions. Financial penalties and technology bans, while effective enforcement tools, underscore the necessity of scrupulous algorithmic investigations (DOJ 2024; Milstein 2024; FTC 2024). These enforcement patterns inform our identified research opportunities, particularly around developing standardized disparate impact measurements and techniques for finding equally accurate, less discriminatory alternatives.
@@
Under US civil rights law, defendants in discrimination cases can justify practices with disparate impact if they demonstrate “business necessity.” However, this defense fails if plaintiffs can show the existence of less discriminatory alternatives that achieve similar business objectives. This concept has direct relevance to algorithmic discrimination cases. In 2022, a class action lawsuit of homeowners alleged that State Farm’s homeowner insurance claims processing algorithms subjected Black policyholders to greater scrutiny. The complaint specifically questioned, “Whether substantially equally or more valid alternative means of claim processing are available that would eliminate or reduce the discriminatory impact [of State Farm’s claims processing policy]” (Huskey 2022). As of writing, the investigation remains ongoing as the plaintiffs have demonstrated a probable likelihood that their allegations have merit (United States District Court 2023). In 2007, a similar case emerged when plaintiffs sued Allstate for allegedly charging racially discriminatory insurance premiums. While denying these claims, Allstate agreed to implement a new pricing algorithm as part of the settlement (DeHoyos 2007). Though the settlement documents do not describe the development of this alternative algorithm, it seems reasonable to assume that it was designed to balance non-discrimination with business performance requirements. Legal scholars have established that defendants have an affirmative duty to search for less discriminatory alternatives in many contexts (Black et al. 2024), and recent work has begun operationalizing this search in fair lending contexts (Gillis, Meursault, and Ustun 2024).
@@
In 2021, the New York State Department of Financial Services (NYDFS) investigated Apple based on consumer complaints about the Apple Card and the potential for algorithmic discrimination. After a lengthy investigation, the report found no evidence of unlawful discrimination against applicants under fair lending law. Part of the investigation focused on individuals who sent discrimination complaints to the NYDFS: “In each instance, the Bank was able to identify the factors that led to the credit decisions, such as credit score, indebtedness, income, credit utilization, missed payments, and other credit history elements” (New York State Department of Financial Services 2021b). Using regression analysis, investigators were able to determine that the Apple Card lending policy—and the underlying statistical model—would not produce disparate impacts (New York State Department of Financial Services 2021b).
@@
Our analysis suggests several critical areas for future research in quantifying algorithmic disparate impact. As algorithmic systems become more complex, they require sophisticated measurement approaches that can: (1) isolate interaction effects between protected attributes and other variables, (2) track temporal drift in model behavior and fairness metrics, and (3) model complex downstream economic consequences. Machine learning concepts such as distribution shift, causal inference, and feature attribution are crucial for these research questions. However, more work is needed to develop practical solutions. For example, in lending discrimination, impact quantification must account for both immediate effects of loan denials and long-term consequences for wealth accumulation and inter-generational mobility. Methods from econometrics and causal inference may help address this change. While prior works advocate for causal frameworks in measuring disparate impact (Xiang and Raji 2019; Xiang 2020), the field lacks validated methodologies and case studies demonstrating how to compute these measures in ways that both withstand legal scrutiny and support specific damage calculations. The few examples that do exist rely on quasi-experimental methods that are often infeasible to implement with the available data (Arnold, Dobbie, and Hull 2021). Recent work has highlighted that regulatory approaches must consider threshold variations across different populations (Meursault et al. 2025) and has begun developing frameworks for explainable fairness in regulatory auditing contexts (O’Neil, Sargeant, and Appel 2024). This methodological gap significantly hampers enforcement actions seeking concrete remedies for affected populations.
@@
Without accurate demographic information, investigators may struggle to establish statistical evidence of discriminatory patterns and face heightened burdens of proof in legal proceedings. The lack of reliable race and ethnicity data particularly hampers class certification efforts, as courts require clear methods for identifying affected individuals. As an extreme case, certain sectors may forbid the consideration—and therefore collection—of these groups, such as the Equal Credit Reporting Act, which prohibits creditors from considering information about protected attributes in any aspect of a credit transaction (CFPB 2011; Consumer Financial Protection Bureau 2018). The lack of recorded information affected the Allstate 2007 settlement regarding allegedly racially discriminatory insurance premiums because “Allstate does not maintain information on the race of its policyholders,” and therefore notice could not be given to prospective class members (DeHoyos 2007).